### v3

# Feature learning as alignment: a structural property of gradient descent in non-linear neural networks

Abstract
Understanding the mechanisms through which neural networks extract statistics from input-label pairs through feature learning is one of the most important unsolved problems in supervised learning. Prior works demonstrated that the gram matrices of the weights (the neural feature matrices, NFM) and the average gradient outer products (AGOP) become correlated during training, in a statement known as the neural feature ansatz (NFA). Through the NFA, the authors introduce mapping with the AGOP as a general mechanism for neural feature learning. However, these works do not provide a theoretical explanation for this correlation or its origins. In this work, we further clarify the nature of this correlation, and explain its emergence. We show that this correlation is equivalent to alignment between the left singular structure of the weight matrices and the newly defined pre-activation tangent features at each layer. We further establish that the alignment is driven by the interaction of weight changes induced by SGD with the pre-activation features, and analyze the resulting dynamics analytically at early times in terms of simple statistics of the inputs and labels. Finally, motivated by the observation that the NFA is driven by this centered correlation, we introduce a simple optimization rule that dramatically increases the NFA correlations at any given layer and improves the quality of features learned.
**footnotetext: Correspondence to: Daniel Beaglehole (dbeaglehole@ucsd.edu)$\dagger$\$$\dagger$\$footnotetext: Work partially done as an intern at Google DeepMind.
1 Introduction
Neural networks have emerged as the state-of-the-art machine learning methods for seemingly complex tasks, such as language generation [1], image classification [2], and visual rendering [3]. The precise reasons why neural networks generalize well have been the subject of intensive exploration, beginning with the observation that standard generalization bounds from statistical learning theory fall short of explaining their performance [4].
A promising line of work emerged in the form of the neural tangent kernel, connecting neural networks to kernels in the wide limit [5, 6]. However, subsequent research showed that the success of neural networks relies critically on aspects of learning which are absent in kernel approximations [7, 8, 9, 10, 11]. Other work showed that low width suffices for gradient descent to achieve arbitrarily small test error [12], further refuting the idea that extremely wide networks are necessary.
Subsequently, the success of neural networks has been largely attributed to feature learning - the ability of neural networks to learn statistics, measurements, and representations of data which are useful for downstream tasks. However, the specific mechanism through which features are learned is an important unsolved problem in deep learning theory. A number of works have studied the abilities of neural networks to learn features in structured settings [13, 14, 15, 16, 17, 18, 19]. Some of that work proves strict separation in terms of sample complexity between neural networks trained with stochastic gradient descent and kernels [20].
The work above studies simple structure, such as learning from low-rank data or functions that are hierarchical compositions of simple elements. Recent work makes a big step towards generalizing these assumptions by proposing the neural feature ansatz (NFA) [21, 22], a general structure that emerges in the weights of trained neural networks. The NFA states that the gram matrix of the weights at a given layer (known as the neural feature matrix or NFM) is aligned with the average gradient outer product (AGOP) of the network with respect to the input to that layer. In particular, the NFM and AGOP are highly correlated in all layers of trained neural networks of general architectures, including practical models such as VGG [23], vision transformers [24], and GPT-family models [1].
A major missing element is an explanation for how and why the AGOP and NFM become correlated through training with gradient descent. In this paper, we precisely describe the emergence of this correlation. We establish that the NFA is equivalent to alignment between the left singular structure of the weight matrices and the uncentered covariance of the pre-activation tangent kernel (PTK) (Section 2) features. We then introduce the centered neural feature correlation (C-NFC) which isolates this alignment process. We show empirically that the C-NFC is close to its maximum value of $1$ at early times, and fully captures the NFA at late times for a variety of architectures (fully-connected, convolutional, and attention layers) over a diverse collection of datasets (Section 3). Our experiments suggest that the C-NFC drives the development of the NFA. Through this centering, we identify analytically how the NFA emerges as a structural property of gradient descent. We provide further evidence by showing that the dynamics of the C-NFC can be understood analytically in terms of the statistics of the data and labels at early times (Section 3), and can be manipulated theoretically (Section 4).
Finally, motivated by our theory, we design an intervention to increase the influence of the C-NFC and make the NFA more robust: Speed Limited Optimization, a layerwise gradient normalization scheme (Section 5). The effectiveness of the latter update rule suggests a path towards rational design of architectures and training procedures that maximize the NFA notion of feature learning by promoting alignment dynamics.
2 Alignment between the weight matrices and the pre-activation tangents
In this section, we decompose the AGOP into the weight matrices and the feature covariance of the pre-activation tangent kernel (PTK), and demonstrate that the NFA is equivalent to alignment between the weights and these PTK features. We include a glossary of terms for ease of reference in Appendix A.
2.1 Preliminaries
We consider fully-connected neural networks with a single output of depth $L\geq 1$, where $L$ is the number of hidden layers, written $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. We write the input to layer $\ell\in\{0,\ldots,L\}$ as $x_{\ell}$, where $x_{0}\equiv x$ is the original datapoint, and the pre-activation as $h_{\ell}(x)$. Then,
$h_{\ell}(x)=W^{(\ell)}x_{\ell},\qquad x_{\ell+1}=\phi(h_{\ell}(x))~{},$ (1)
where $\phi$ is an element-wise nonlinearity, $W^{(\ell)}\in\mathbb{R}^{k_{\ell+1}\times k_{\ell}}$ is a weight matrix, and $k_{\ell}$ is the hidden dimension at layer $\ell$. We restrict $k_{L+1}$ to be the number of output logits, and set $k_{0}=d$, where $d$ is the input dimension of the data. Note that $f(x)=h_{L}(x)$ and we ignore $x_{L+1}$. We train $f$ by gradient descent on a loss function $\mathcal{L}(\theta,X)$, where $X$ is an input dataset, and $\theta$ is the collection of weights.
We consider a supervised learning setup where we are provided $n$ input-label pairs $(x^{(1)},y^{(1)}),\ldots,(x^{(n)},y^{(n)})\in\mathbb{R}^{d}\times\mathbb{R}$. We denote the network inputs (datapoints) $X\in\mathbb{R}^{n\times d}$ and the labels $y\in\mathbb{R}^{n\times 1}$. For a given network, the inputs to intermediate layers $\ell\in\{0,\ldots,L\}$ are written $X_{\ell}\in\mathbb{R}^{n\times k_{\ell}}$, where $X_{0}\equiv X$. We train a fully-connected neural network to learn the mapping from network inputs to labels by minimizing a standard loss function, such as mean-squared-error or cross-entropy, on the dataset.
Figure 1: Various feature learning measures for target function $y(x)=\sum_{k=1}^{r}x_{k\text{ mod }r}\cdot x_{(k+1)\text{ mod }r}$ with $r=5$ and inputs drawn from standard normal. The EGOP $\mathbb{E}_{x\sim\mu}\left[\frac{\partial y}{\partial x}\frac{\partial y}{\partial x}^{\top}\right]$ (first plot) captures the low-rank structure of the task. The NFM $\left(W^{\top}W\right)$ (second plot) and AGOP $\left(W^{\top}KW\right)$ (third plot) of a fully-connected network are similar to each other and the EGOP. Replacing $K$ with a symmetric matrix $Q$ with the same spectrum but independent eigenvectors obscures the low rank structure (fourth plot), and reduces the correlation from $\rho\left(F,\bar{G}\right)=0.93$ to $\rho\left(F,W^{\top}QW\right)=0.53$.
One can define two objects associated with neural networks that capture learned structure. For a given layer $\ell$, the neural feature matrix (NFM) $F_{\ell}$ is the gram matrix of the columns of the weight matrix $W^{(\ell)}$, i.e. $F_{\ell}\equiv(W^{(\ell)})^{\top}W^{(\ell)}$. $F_{\ell}$ depends on the right singular vectors (and corresponding singular values) of $W^{(\ell)}$. The second fundamental object we consider is the average gradient outer product (AGOP) $\bar{G}_{\ell}$, defined as $\bar{G}_{\ell}\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial x_{\ell}}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial x_{\ell}}^{\top}$.
Since both these matrices have the same dimensions ($k_{\ell+1}\times k_{\ell+1}$), we can consider their cosine similarity. We define the neural feature correlation (NFC) by:
$\rho\left(F_{\ell},\bar{G}_{\ell}\right)\equiv\text{tr}\left(F_{\ell}^{\top}\bar{G}_{\ell}\right)\cdot\text{tr}\left(F_{\ell}^{\top}F_{\ell}\right)^{-1/2}\cdot\text{tr}\left(\bar{G}_{\ell}^{\top}\bar{G}_{\ell}\right)^{-1/2}~{}.$ (NFC)
This takes on values in $[0,1]$ since $F_{\ell}$ and $\bar{G}_{\ell}$ are both PSD.
Prior work has shown that in trained neural networks, the NFC will generally be close to $1$ to varying degree of approximation. This notion is formalized in the Neural Feature Ansatz (NFA):
Ansatz 1 (Neural Feature Ansatz [21]).
The Neural Feature Ansatz states that, for all layers $\ell\in[L]$ of a fully-connected neural network with $L$ hidden layers trained on input data $x^{(1)},\ldots,x^{(n)}$, the NFC will satisfy $\rho\left(\bar{G}_{\ell},F_{\ell}\right)\approx 1~{}.$
The inputs to the covariance are the NFM and AGOP with respect to the input of layer $\ell$. Here, $\frac{\partial f(x)}{\partial x_{\ell}}\in\mathbb{R}^{k_{\ell}\times 1}$ denotes the gradient of the function $f$ with respect to the intermediate representation $x_{\ell}$. For simplicity, we may concatenate these gradients across the input dataset into a single matrix $\frac{\partial f(X)}{\partial x_{\ell}}\in\mathbb{R}^{n\times k_{\ell}}$. Note we consider scalar outputs in this work, though the NFA relation is identical when there are $c\geq 1$ outputs, where in this case $\frac{\partial f(x)}{\partial x_{\ell}}\in\mathbb{R}^{k_{\ell}\times c}$ is the input-output Jacobian of the model $f$.
We note that while the ansatz states exact proportionality, in practice, the NFM and AGOP are highly correlated with correlation less than $1$, where the final correlation depends on many aspects of training and architecture choice (discussed further in Section 3).
The relation between the NFM and the AGOP is significant, in part, because for a neural network that has learned enough information about the target function, the AGOP of this model with respect to the first-layer inputs will approximate the expected gradient outer product (EGOP) of the target function [25]. In particular, the EGOP of the target function contains task-specific structure that is completely independent of the model used to estimate it. Where the labels are generated from a particular target function $y(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ on data sampled from a distribution $\mu$, the EGOP is defined as
$\operatorname{EGOP}(y,\mu)=\mathbb{E}_{x\sim\mu}\left[\frac{\partial y}{\partial x}\frac{\partial y}{\partial x}^{\top}\right]~{}.$
If the neural feature ansatz (Ansatz 1) holds, the correlation of the EGOP and the AGOP at the end of training also implies high correlation between the NFM of the first layer and the EGOP, so that the NFM has encoded this task-specific structure.
To demonstrate the significance of high correlation between the NFM and the AGOP in successfully trained networks, we consider the following chain-monomial low-rank task:
$y(x)=\sum_{k=1}^{r}x_{k\text{ mod }r}\cdot x_{(k+1)\text{ mod }r}~{},$ (2)
where the data inputs are sampled from an isotropic Gaussian distribution $\mu=\mathcal{N}(0,I)$. In this case, the entries $\operatorname{EGOP}(y,\mu)$ will be 00 for rows and columns outside of the $r\times r$ sub-matrix corresponding to $x_{1},\ldots,x_{r}$ (Figure 1), as $y$ does not vary with coordinates $x_{r+1},\ldots,x_{d}$. Within this sub-matrix, the diagonal entries will have value $2$, while the off-diagonal entries will be either $1$ or 00. Therefore, $\operatorname{EGOP}(y,\mu)$ will be rank $r$, where $r$ is much less than the ambient dimension.
We verify for this task that the AGOP of the trained model resembles the EGOP (first and third panels of Figure 1). Here the NFA holds and the NFM (second panel) resembles the AGOP and therefore the EGOP as well. Therefore, the neural network has learned the model-independent and task-specific structure of the chain-monomial task in the right singular values and vectors of the first layer weight matrix, as these are determined by the NFM. In fact, previous works have demonstrated that the NFM of the first layer of a well-trained neural network is highly correlated with the AGOP of a fixed kernel method trained on the same dataset [21].
This insight has inspired iterative kernel methods which can match the performance of fully-connected networks [21, 26, 27] and improve over fixed convolutional kernels [22]. Additional prior works demonstrate the benefit of including the AGOP features to improve feature-less predictors [28, 29, 30]. The AGOP has also been show to capture surprising phenomena of neural networks beyond low-rank feature learning including deep neural collapse [31].
2.2 Alignment decomposition
In order to understand Ansatz 1, it is useful to decompose the AGOP. Doing so will allow us to show that the neural feature correlation (NFC) can be interpreted as an alignment metric between weight matrices and the pre-activation tangent kernel (PTK). The PTK $\mathcal{K}^{(\ell)}(x,z)$ is defined with respect to a layer $\ell$ of a neural network and two inputs $x,z$. The kernel evaluates to:
$\mathcal{K}^{(\ell)}(x,z)\equiv\frac{\partial f(x)}{\partial h_{\ell}}\cdot\frac{\partial f(z)}{\partial h_{\ell}}~{}.$ (PTK)
When the arguments $x$ and $z$ are omitted, $\mathcal{K}^{(\ell)}\in\mathbb{R}^{n\times n}$ consists of the matrix of kernel evaluations on all pairs of datapoints. We may also consider the covariance $K^{(\ell)}\in\mathbb{R}^{k_{\ell}\times k_{\ell}}$ of the features associated with the PTK, $\frac{\partial f(X)}{\partial h_{\ell}}\in\mathbb{R}^{n\times k_{\ell}}$. In particular, we define the feature covariance of the PTK as,
$K^{(\ell)}\equiv\frac{\partial f(X)}{\partial h_{\ell}}^{\top}\frac{\partial f(X)}{\partial h_{\ell}}~{}.$ (PTK feature covariance)
Crucially, for any layer $\ell$, we can re-write the AGOP in terms of this feature covariance as,
$\bar{G}_{\ell}=(W^{(\ell)})^{\top}K^{(\ell)}W^{(\ell)}~{}.$
This gives us the following proposition:
Proposition 2 (Alignment decomposition of NFC).
$\rho\left(F_{\ell},\bar{G}_{\ell}\right)=\rho\left((W^{(\ell)})^{\top}W^{(\ell)},(W^{(\ell)})^{\top}K^{(\ell)}W^{(\ell)}\right)~{}.$
This alignment holds trivially and exactly if $K^{(\ell)}$ is the identity. However, the correlation can be high in trained networks where $K^{(\ell)}$ is non-trivial. For example, in the chain monomial task (Figure 1), $K^{(0)}$ is far from identity (standard deviation of its eigenvalues is $5.9$ times its average eigenvalue), but the NFA correlation is $0.93$ at the end of training. We also note that if $K^{(\ell)}$ is independent of $W^{(\ell)}$, the alignment is lower than in trained networks; in the same example, replacing $K^{(0)}$ with a matrix $Q$ with equal spectrum but random eigenvectors greatly reduces the correlation to $0.53$ and qualitatively disrupts the structure relative to the NFM (Figure 1, rightmost column). We show the same result for the CelebA dataset (see Appendix L). Therefore, the NFA is a consequence of alignment between the left eigenvectors of $W^{(\ell)}$ and $K^{(\ell)}$ in addition to spectral considerations.
Note that $K^{(\ell)}$ itself is the AGOP of the neural network $f$ with respect to the pre-activations at layer $\ell$, i.e. $K^{(\ell)}\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial h_{\ell}}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial h_{\ell}}^{\top}$. For convolutional and attention layers, the $K^{(\ell)}$ are computed by additionally averaged over all patches and token positions in the input, respectively (see Radhakrishnan et al. [21], Beaglehole et al. [22] for the formulation of the NFA in these architectures).
Figure 2: Uncentered and centered neural feature correlations across (A,B) fully-connected, (C) convolutional, and (D) attention layers with large initialization scale. (A,C,D) show trajectories of C/UC-NFC over training. (B) shows NFC values across all layers of an MLP with five hidden layers, averaged over CIFAR-10, CIFAR-100, SVHN, MNIST, GTSRB, and STL-10 datasets. (A-C) are additionally averaged over three random seeds. Each row of (D) is an attention block (ordered from first to last in the GPT model), while the columns show correlations for query, key, and value layers, respectively.
3 Centering the NFC isolates weight-PTK alignment
We showed that the neural feature ansatz is equivalent to PTK-weight alignment (Proposition 2). We now ask: is the increase in the NFC due to alignment of the weight matrices to the current PTK, or the alignment of the PTK to the current weights? In practice, both effects matter, but numerical evidence suggests that changes in the PTK do not drive the early dynamics of the NFC (Appendix C). Instead, we observe that the alignment between the weights and the PTK feature covariance is driven by a centered NFC that captures alignment between the parameter changes and the PTK. We then show this centered NFC can hold robustly in settings where the NFC holds with correlation less than 1, such as early in training and/or with large initialization. Finally, we establish analytically through the centered NFC how the NFA emerges as a structural property of gradient descent.
We begin by considering a decomposition of the NFM and AGOP into parts that depend on initialization and a part that depends on the changes in the weight matrix. Let $W^{(\ell)}_{t}$ and $K^{(\ell)}_{t}$ be the weight matrix and PTK feature covariance for layer $\ell$ at time $t$. We can write the NFM and AGOP in terms of the initial weights $W^{(\ell)}_{0}$ and the change in weights $\bar{W}^{(\ell)}_{t}\equiv W^{(\ell)}_{t}-W^{(\ell)}_{0}$ as follows:
$\begin{split}F{}=W_{t}^{\top}W_{t}&=\bar{W}_{t}^{\top}\bar{W}_{t}+W_{0}^{\top}\bar{W}_{t}+\bar{W}_{t}^{\top}W_{0}+W_{0}^{\top}W_{0},\\ \bar{G}{}=W_{t}^{\top}K_{t}W_{t}&=\bar{W}_{t}^{\top}K_{t}\bar{W}_{t}+W_{0}^{\top}K_{t}\bar{W}_{t}+\bar{W}_{t}^{\top}K_{t}W_{0}+W_{0}^{\top}K_{t}W_{0}\end{split}$ (3)
where we omitted $\ell$ from all terms for ease of notation. The first term in each decomposition isolates the gram matrix of the changes in the weight matrix. We call $\bar{W}^{\top}\bar{W}$ the centered NFM and $\bar{W}^{\top}K\bar{W}$ the centered AGOP. The centered AGOP in particular measures the alignment of weight updates with the current PTK feature covariance. Both terms are 00 at initialization, and if the weight matrices change significantly (that is, if $||\bar{W}||\gg||W_{0}||$), both the NFM and AGOP are dominated by the centered terms.
The decomposition in Equation 3 suggests that updates which drive correlation of the centered NFM and centered AGOP can also drive the overall value of the original NFC. Inspired by this observation, we define the centered NFC (C-NFC) as,
$\rho\left((\bar{W}^{(\ell)})^{\top}\bar{W}^{(\ell)},(\bar{W}^{(\ell)})^{\top}K^{(\ell)}\bar{W}^{(\ell)}\right)~{}.$
For the remainder of the paper, we will refer to the original NFC as the uncentered NFC (UC-NFC) to avoid ambiguity.
The centered NFC is consistently higher than the uncentered NFC across training times, architectures, and datasets Figure 2. This especially holds at early times and in deeper layers of a network (Figure 2C and D, VGG-11 on CIFAR-10 and the GPT model on Shakespeare respectively). For MLPs, we conduct a broader set of experiments and verify that the trends hold on a wide range of vision datasets. We also note that the C-NFC at early times is relatively robust to the initialization statistics of the weight matrix $W^{(\ell)}$ unlike the UC-NFC (for additional experiments, see Appendix G). We will return to this point in Section 5. These experiments suggest the C-NFC is responsible for improvements in the uncentered correlation.
High C-NFC values can drive increases in the UC-NFC as long as the centered NFM and centered AGOP are increasing in magnitude during training. We can confirm that the weights move significantly from initialization, which drives the contribution from the centered NFM and centered AGOP (Figures 14, and 15 in Appendix L). The increased importance of the C-NFC leads the UC-NFC to converge to the C-NFC at late times in most of our experimental settings, as the contribution from the weight changes dominates the contribution from the initialization. These findings validate that the C-NFC is an important contributor to increases in the UC-NFC.
Our experiments suggest that studying the C-NFC is a useful first step to understanding the neural feature ansatz. In the remainder of this section, we develop a theoretical analysis to understand why the C-NFC is generically large at early times. We then use this analysis to motivate interventions which can keep the C-NFC high and promote the NFA earlier on in training (Section 5).
3.1 Analysis at early times
We now theoretically identify why the centered AGOP and NFM become correlated as a structural property of gradient descent, for at least the early training times. We consider the setup introduced in Equation (1). For theoretical convenience we focus on the case of training under gradient flow, where the dynamics of a weight matrix $W^{(\ell)}$ trained on loss $\mathcal{L}$ is given by
$\dot{W}^{(\ell)}\equiv\frac{dW^{(\ell)}}{dt}=-\nabla_{W^{(\ell)}}\mathcal{L}~{},$ (4)
where $\dot{W}^{(\ell)}$ is the time derivative of $W^{(\ell)}$. We expect our results to apply to gradient descent with small and intermediate learning rates as well. For simplicity of notation, we omit the layer index $\ell$.
We note that $\bar{W}=0$ at initialization. This gives us the following proposition, which shows that early time dynamics of the centered NFM and AGOP are dominated by the second time derivatives:
Proposition 3 (Centered NFC dynamics).
Let $W$ be the weights of a fully-connected layer of a neural network at initialization and $X$ be the inputs to that layer. Then, when the neural network is trained by gradient flow on a loss function $\mathcal{L}$, we have $\bar{W}^{\top}\bar{W}=\bar{W}^{\top}K\bar{W}=0$, $\frac{\operatorname{d}}{\operatorname{d}t}(\bar{W}^{\top}\bar{W})=\frac{\operatorname{d}}{\operatorname{d}t}(\bar{W}^{\top}K\bar{W})=0$, and the first non-trivial time derivatives satisfy,
$\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}\bar{W})=2\cdot X^{\top}\dot{\mathcal{L}}\mathcal{K}\dot{\mathcal{L}}X,\quad\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}K\bar{W})=2\cdot X^{\top}\dot{\mathcal{L}}\mathcal{K}^{2}\dot{\mathcal{L}}X~{}$ (5)
where $\dot{\mathcal{L}}$ is the $n\times n$ diagonal matrix of logit derivatives ˙$\dot{\mathcal{L}}\equiv{{\rm diag}\left(\frac{\partial\mathcal{L}}{\partial f}\right)}$.
Proof of Proposition 3.
At initialization, all terms containing at least one copy of $\bar{W}$ vanish, leading to the zero and first time derivatives to be 0. Further, we have $\dot{W}=\frac{\partial f(X)}{\partial h}^{\top}\dot{\mathcal{L}}X$. Therefore, using that $K\equiv\frac{\partial f(X)}{\partial h}^{\top}\frac{\partial f(X)}{\partial h}$ and $\mathcal{K}\equiv\frac{\partial f(X)}{\partial h}\frac{\partial f(X)}{\partial h}^{\top}$,
$\frac{1}{2}\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}\bar{W})=\dot{W}^{\top}\dot{W}=X^{\top}\dot{\mathcal{L}}\frac{\partial f(X)}{\partial h}\frac{\partial f(X)}{\partial h}^{\top}\dot{\mathcal{L}}X=X^{\top}\dot{\mathcal{L}}\mathcal{K}\dot{\mathcal{L}}X~{},$
and,
$\frac{1}{2}\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}\bar{W})=\dot{W}^{\top}K\dot{W}$ $=X^{\top}\dot{\mathcal{L}}\frac{\partial f(X)}{\partial h}\,K\,\frac{\partial f(X)}{\partial h}^{\top}\dot{\mathcal{L}}X$
$=X^{\top}\dot{\mathcal{L}}\frac{\partial f(X)}{\partial h}\frac{\partial f(X)}{\partial h}^{\top}\frac{\partial f(X)}{\partial h}\frac{\partial f(X)}{\partial h}^{\top}\dot{\mathcal{L}}X$
$=X^{\top}\dot{\mathcal{L}}\mathcal{K}^{2}\dot{\mathcal{L}}X~{}.$
We immediately see from these equations how gradient-based training can drive the C-NFC to large (close to $1$) initial values - the first non-trivial time derivatives are highly correlated. This also sheds light on another mystery: namely, that the NFC does not explicitly involve the loss, yet it holds after gradient based training on the loss. The C-NFC is driven by the fact that the dynamics of the centered NFM and AGOP depend on the labels in a similar way, differing only by a single factor of the kernel $\mathcal{K}$. This eventually drives the UC-NFC and causes the NFA to hold.
When $\mathcal{K}$ is proportional to a projection matrix, i.e. $\mathcal{K}\propto\mathcal{K}^{2}$, these derivatives will be perfectly correlated. Even in the general case where $\mathcal{K}$ is dispersed with moderate eigenvalue decay, and does not commute with ˙$\dot{\mathcal{L}}XX^{\top}\dot{\mathcal{L}}$, the two quantities have non-negative covariance. In the next section we will construct a dataset which interpolates between adversarial and aligned eigenstructure to demonstrate the range of possible values the derivative correlations can take.
In the remainder of this work, we first analyze the C-NFC in early training and show that gradient based training drives the C-NFC, with a strength that depends on model and datasets (Section 4). We then discuss initializations and optimizers which can increase the C-NFC and induce larger correlation between $F$ and $\bar{G}$ (Section 5).
4 Theoretically predicting the centered NFC
In this section, we show how our decomposition for the time derivatives of the centered NFC in early times can be used to predict its value theoretically.
4.1 Early time C-NFC dynamics
In order to understand the correlation of centered AGOP and NFM more quantitatively, we will focus on the case of mean-squared error (MSE) loss. Here $\dot{\mathcal{L}}$ corresponds to the diagonal matrix of the residuals $y-f(x)$. If the outputs of the network is 00 on the training data, then ˙$\dot{\mathcal{L}}=Y\equiv{\rm diag}\left(y\right)$, the labels themselves. In that case, the correlation of the time derivatives at initialization is a function of the input-label covariances as well as the PTK matrix:
$\begin{split}&\rho\left(\dot{W}^{\top}\dot{W},\dot{W}^{\top}K\dot{W}\right)=\,\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)\\ &\cdot\text{tr}\left((X^{\top}Y\mathcal{K}YX)^{2}\right)^{-1/2}\cdot\text{tr}\left((X^{\top}Y\mathcal{K}^{2}YX)^{2}\right)^{-1/2}\end{split}$ (6)
In general we expect that these quantities can self-average under the appropriate high-dimensional limits. It is helpful then to write down the average (over initializations) of the covariance. Taking expectation of the first factor,
$\begin{split}&\mathbb{E}_{\theta}\left[\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)\right]=\\ &\text{tr}\left(X^{\top}Y\mathbb{E}_{\theta}\left[\mathcal{K}\right]YXX^{\top}Y\mathbb{E}_{\theta}\left[\mathcal{K}^{2}\right]YX\right)\\ &+\text{tr}\left(\text{Cov}\left(X^{\top}Y\mathcal{K}YX,X^{\top}Y\mathcal{K}^{2}YX\right)\right)\end{split}$ (7)
A similar decomposition exists for the denominator terms.
The simplest limit is the NTK regime where the width tends to infinity with fixed number of data points and input dimension. Here the PTK matrix will approach its expectation, and the second, covariance term above will tend to 0. This suggests that the first (mean) term encodes much of the interesting phenomenology near initialization for large networks. We use this fact to design a dataset that interpolates between small and large initial derivatives of the C-NFC, even in the non-NTK regime, which will help us demonstrate the validity of our random matrix theory approach.
We focus most of our random matrix theory analysis on a one-hidden layer network, as the calculations quickly become complicated with depth. However, we also show that the mean term can be used to approximately predict the NFC value for Gaussian data with different spectral decay rates for more complicated networks (see Appendix D).
4.2 Exact predictions with one hidden layer and quadratic activations
Figure 3: Predicted versus observed correlation of the second derivatives of centered $F$ and $\bar{G}$ on the alignment reversing dataset. Different shaded color curves correspond to four different seeds for the dataset. The solid blue curve is the average over all data seeds. The rightmost sub-figure is a scatter plot of the predicted versus observed correlations of these second derivatives, with one point for each balance value. We instantiate the dataset in the proportional regime where width, input dimension, and dataset size are all equal to $1024$.
We can capture the behavior of both the mean and covariance terms from Equation (7) in certain high dimensional settings. In particular for one hidden layer neural networks with quadratic activations, we can exactly predict the value of Equation (7) in the high dimensional limit. We make the assumption that $X$ and $Y$ are (asymptotically) freely independent of the parameters at initialization, and that the resulting average is close to the value of any specific network initialization (self-averaging). For more details on the assumptions, see Appendix E. We then show numerically that the predicted correlation between the second derivatives of the centered NFM and AGOP matches the observed correlation across individual random instances of the dataset.
To illustrate our analysis, we derive here the numerator of Equation (6). With one hidden-layer, the neural network simplifies to $f(x)=a^{\top}(Wx)^{2}$, where $a$ is the readout layer, and the square is element-wise. Let $M^{(4)}_{X|Y}=(X^{\top}YX)^{2}$, and $M^{(2)}_{X}=X^{\top}X$, and $F_{a}=W^{\top}{\rm diag}\left(a^{2}\right)W$. Then,
$\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)=\text{tr}\left(M^{(4)}_{X|Y}F_{a}M^{(4)}_{X|Y}F_{a}M^{(2)}_{X}F_{a}\right)~{}.$
We can simplify this alternating product using standard results from random matrix theory, and factor the parameter and data contributions at the cost of increasing the number of terms in the expression (Appendix E).
In the simplified case of isotropic data $X^{\top}X=I$, we have:
$\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)=\text{tr}\left(M^{(4)}_{X|Y}\right)^{2}\text{tr}\left(F_{a}^{3}\right)+\text{tr}\left(F_{a}\right)\text{tr}\left(F_{a}^{2}\right)\text{tr}\left((\bar{M}^{(4)}_{X|Y})^{2}\right)~{}.$
From these random matrix calculations, we see that the correlations of the derivatives are determined by traces of powers of $M^{(4)}_{X|Y}$ (and $M^{(2)}_{X}$ by the calculations in Section E), which is specific to the dataset, and $F_{a}$, which is specific to the architecture and initialization.
Manipulating the C-NFC
To numerically explore the validity of the random matrix theory calculations, we need a way to generate datasets with different values of $\rho\left(\dot{W}^{\top}\dot{W},\dot{W}^{\top}K\dot{W}\right)$. We construct a random dataset called the alignment reversing dataset, parameterized by a balance parameter $\gamma\in(0,1]$, to adversarially disrupt the NFA near initialization in the regime that width $k$, input dimension $d$, and dataset size $n$ are all equal ($n=k=d=1024$). By Proposition 8, for the aforementioned neural architecture, the expected second derivative of the centered NFM satisfies, $\mathbb{E}\left[\dot{W}^{\top}\dot{W}\right]=X^{\top}Y\mathbb{E}\left[\mathcal{K}\right]YX=(X^{\top}YX)^{2}$, while the expected second derivative of the centered AGOP, $\mathbb{E}\left[\dot{W}^{\top}K\dot{W}\right]=X^{\top}Y\mathbb{E}\left[\mathcal{K}^{2}\right]YX$, has an additional component $X^{\top}YX\cdot X^{\top}X\cdot X^{\top}YX$. Our construction exploits this difference in that $X^{\top}X$ becomes adversarially unaligned to $X^{\top}YX$ as the balance parameter decreases.
The construction exploits that we can manipulate $X^{\top}YX\cdot X^{\top}X\cdot X^{\top}YX$ freely of the NFM using a certain choice of $Y$. We design the dataset such that this AGOP-unique term is close to identity, while the NFM second derivative has many large off-diagonal entries, leading to low correlation between the second derivatives of the NFM and AGOP.
Dataset Construction
The data consists of a mixture of two distributions from which two subsets of the data $X^{(1)}$ and $X^{(2)}$ are sampled from, and is parametrized by the balance $\gamma\in(0,1]$ and two variance parameters $\epsilon_{1},\epsilon_{2}>0$. The subset $X^{(1)}$ which has label $y_{1}=1$ and constitutes a $\gamma$ fraction of the entire dataset, is sampled from a multivariate Gaussian distribution with mean 00 and covariance
$\Sigma=\bm{1}\bm{1}^{\top}+\epsilon_{1}\cdot I~{}.$
Then the second subset, $X^{(2)}$, is constructed such that $(X^{(2)})^{\top}X^{(2)}\approx((X^{(1)})^{\top}X^{(1)})^{-2}$, and has labels $y_{2}=0$. Then, for balance parameter $\gamma$ sufficiently small, the AGOP second derivative approximately satisfies,
$\mathbb{E}\left[\dot{W}^{\top}K\dot{W}\right]$ $\sim X^{\top}YXX^{\top}XX^{\top}YX$
$=(X^{(1)})^{\top}X^{(1)}X^{\top}X(X^{(1)})^{\top}X^{(1)}$
$\approx(X^{(1)})^{\top}X^{(1)}(X^{(2)})^{\top}X^{(2)}(X^{(1)})^{\top}X^{(1)}$
$\approx I~{},$
In contrast, the NFM second derivative, $\mathbb{E}\left[\dot{W}^{\top}\dot{W}\right]=(X^{\top}YX)^{2}=((X^{(1)})^{\top}X^{(1)})^{2}\approx\Sigma^{2}$, will be significantly far from identity.
In our experiment, we sample multiple random datasets with this construction and compute the predicted and observed correlation of the second derivatives of the centered NFC at initialization. For specific details on the construction, see Appendix F.
We observe in Figure 3 that the centered NFC predicted with random matrix theory closely matches the observed values, across individual four random seeds and for the average of the correlation across them. Crucially, a single neural network is used across the datasets, confirming the validity of the self-averaging assumption. The variation in the plot across seeds come from randomness in the sample of the data, which cause deviations from the adversarial construction.
5 Increasing the centered contribution to the NFC strengthens feature learning
Our theoretical and experimental work has established that gradient based training leads to alignment of the weight matrices to the PTK feature covariance. This process is driven by the C-NFC. Therefore, one path towards improving the neural feature correlations is to increase the contribution of the C-NFC to the dynamics - as measured, for example, by the centered-to-uncentered, or, C/UC, ratio $\text{tr}\left(\bar{W}^{\top}\bar{W}\bar{W}^{\top}K\bar{W}\right)\text{tr}\left(W^{\top}WW^{\top}KW\right)^{-1}$. When this ratio is large, the C-NFC contributes significantly in magnitude to the UC-NFC, indicating successful feature learning as measured by the uncentered NFC. We will discuss how small initialization promotes feature learning, and design an optimization rule, Speed Limited Optimization, that increases the C/UC ratio and drives the value of the UC-NFC to $1$.
5.1 Feature learning and initialization
One factor that modulates the level of feature learning is the scale of the initialization in each layer. In our experiments training networks with unmodified gradient descent, we observe that the centered NFC will increasingly dominate the uncentered NFC with decreasing initialization (third panel, first row, Figure 4). Further, as the centered NFC increases in contribution to the uncentered quantity, the strength of the UC-NFC, and to a lesser extent, the C-NFC increases (Figure 4). The decreases in correspondence between these quantities is also associated with a decrease in the feature quality of the NFM (Appendix H), for the chain monomial task.
For fully-connected networks with homogeneous activation functions, such as ReLU, and no normalization layers, decreasing initialization scale is equivalent to decreasing the scale of the outputs, since we can write $f(Wx)=a^{-p}f(aWx)$ for any scalar $a$ for any homogenous activation $f$. This in turn is equivalent to increasing the scale of the labels. Therefore, decreasing initialization forces the weights to change more in order to fit the labels, leading to more change in $F$ from its initialization. Conceptually, this aligns with the substantial line of empirical and theoretical evidence that increasing initialization scale or output scaling transitions training between the lazy and feature learning regimes [6, 32, 33, 34].
This relationship suggests that small initialization can be broadly applied to increase the change in F, and the value of the UC-NFC. However, this may not be ideal; for example, if the activation function is differentiable at 00, small initialization leads to a network which is approximately linear. This may lead to low expressivity unless the learning dynamics can increase the weight magnitude.
5.2 Speed Limited Optimization
Figure 4: The effect of SLO on C/UC neural feature correlations and feature learning on the chain monomial task. In the first two rows, we plot the uncentered and centered NFA for the first layer weight matrix as a function of initialization, (A) with standard training and (B) with SLO. We consider a two hidden layer network with ReLU activations, where we set $C_{0}=500$, and $C_{1}=C_{2}=0.002$. The third column shows the ratio of the unnormalized C-NFC to the UC-NFC: ${\text{tr}\left(\bar{W}^{\top}\bar{W}\bar{W}^{\top}K\bar{W}\right)\cdot{\text{tr}\left(W^{\top}WW^{\top}KW\right)}^{-1}}$. The fourth column shows the training loss. In the third row, we plot the NFM and AGOP from a trained network with (C) standard training and (D) with Speed Limited Optimization with fixed initialization scale of $1.0$.
We can instead design an intervention which can increase feature learning without the need to decrease the initialization scale. We do so by fixing the learning speed layerwise to constant values, which causes the C-NFC to dominate the UC-NFC dynamics. For weights at layer $\ell$ and learning rate $\eta>0$, we introduce Speed Limited Optimization (SLO), which is characterized by the following update rule,
$W^{(\ell)}_{t+1}\leftarrow W^{(\ell)}_{t}+\eta\cdot C_{\ell}\cdot\frac{\nabla_{W^{(\ell)}_{t}}\mathcal{L}}{\|\nabla_{W^{(\ell)}_{t}}\mathcal{L}\|}$
where the hyperparameter $C_{\ell}\geq 0$ controls the amount of learning in layer $\ell$. We expect this rule to increase the strength of the UC-NFC in layers where $C_{\ell}$ is large relative to $C_{m}$ for $m\neq\ell$, as $W^{(\ell)}$ will be forced to change significantly from initialization. As a result, $\left\|W_{0}^{(\ell)}\right\|\left\|W_{t}^{(\ell)}\right\|^{-1}\rightarrow 0$ for large $t$, causing the centered and uncentered NFC to coincide for this layer.
We demonstrate the effects of SLO on the chain-monomial task (Equation (2)). We found that fixing the learning speed to be high in the first layer and low in the remaining layers causes the ratio of the unnormalized C-NFC to the UC-NFC to become close to $1$ across initialization scales (Figure 4). The same result holds for the SVHN dataset (see Appendix L). We note that this intervention can be applied to target underperforming layers and improve generalization in deeper networks (see Appendix J for details).
We observe that both the C-NFC and the UC-NFC become close to $1$ after training with SLO, independent of initialization scale (Appendix H). Further, the quality of the features learned, measured by the similarity of $F$ and $\bar{G}$ to the true EGOP, significantly improve and are more similar to each other with SLO, even with large initialization. In contrast, in standard training the UC-NFA fails to develop with large initialization, as $F$ resembles identity (no feature learning), while $\bar{G}$ only slightly captures the relevant features.
Speed Limited Optimization is a step toward the design of optimizers that improve generalization and training times through maximizing feature learning.
6 Discussion
Centered NFC through training
Our theoretical analysis in this work shows that the PTK feature covariance at initialization has a relatively simple structure in terms of the weights and the data. To predict the NFC later in training, we will likely need to account for the change in this matrix. One should be able to predict this development at short times by taking advantage of the fact that eigenvectors of the Hessian change slowly during training [35]. An alternative approach would be to use a quadratic model for neural network dynamics [36, 37] which can capture dynamics of the empirical NTK (and therefore the PTK).
Depth and general activations
We provide an analytic prediction for the C-NFC at initialization for one hidden layer networks with quadratic activation functions, and provide first-order predictions for two hidden layer networks with quadratic and ReLU activations. A followup direction would be to use random matrix analysis to more accurately predict the NFC for deep architectures with general activations. A promising approach in this direction is to approximate non-linear activations with a linear expansion and independent correction terms [38, 39, 40, 41]. Alternatively, we may hope to establish a general spectral condition on $\mathcal{K}$ that facilitates high correlation between $\dot{W}^{\top}\dot{W}$ and $\dot{W}^{\top}K\dot{W}$ at early training times.
Speed Limited Optimization
There are prior works that implement differential learning rates across layers [42, 43], though these differ from differential learning speeds as in SLO, in which the L2-norm of the gradients are fixed, and not the scaling of the gradients. Our intervention is more similar to the LARS [44] and LAMB [45] optimizers that fix learning speeds to decrease training time in ResNet and BERT architectures. However, these optimizers fix learning speeds to the norm of the weight matrices, while our intervention sets learning speeds to free hyperparameters in order to increase feature learning, as measured by the NFC. We also note as an additional explanation for the success of SLO in our experiments, the PTK feature covariance may change slowly (or just in a single direction) with this optimizer, because the learning rate is small for later layers, allowing the first layer weights to align with the PTK.
Adaptive and stochastic optimizers
We demonstrate in this work that the dataset (by our construction in Section 4) and optimizer (with SLO in Section 5) play a significant role in the strength of feature learning. Important future work would be to understand other settings where significant empirical differences exist between optimization choices. In particular, analyzing the NFC may clarify the role of gradient batch size and adaptive gradient methods in generalization [46].
7 Acknowledgements
We thank Lechao Xiao for detailed feedback on the manuscript. We also thank Jeffrey Pennington for helpful discussions. This work used the programs (1) XSEDE (Extreme science and engineering discovery environment) which is supported by NSF grant numbers ACI-1548562, and (2) ACCESS (Advanced cyberinfrastructure coordination ecosystem: services & support) which is supported by NSF grants numbers #2138259, #2138286, #2138307, #2137603, and #2138296. Specifically, we used the NCSA Delta system, via allocations TG-CIS220009.
Appendix A Glossary
  1. 1.
    Expected gradient outer product (EGOP). Defined with respect to a target function $f^{*}$ and an input data distribution $\mu$ over data in $d$ dimensions.
    $\operatorname{EGOP}(f^{*},\mu)\equiv\mathbb{E}_{x\sim\mu}\left[\nabla_{x}f^{*}(x)\nabla_{x}f^{*}(x)^{\top}\right]\in\mathbb{R}^{d\times d}.$
  2. 2.
    Average gradient outer product (AGOP). Defined with respect to a predictor $f$ and a set of inputs $\{x^{(1)},\ldots,x^{(n)}\}$.
    $\mathrm{AGOP}(f,\{x^{(\alpha)}\}_{\alpha})\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)})}{\partial x}\frac{\partial f(x^{(\alpha)})}{\partial x}^{\top}\in\mathbb{R}^{d\times d}.$
    In the context of a deep neural network, we write the AGOP $\bar{G}_{\ell}$ with respect to the inputs of a given layer $\ell$ as
    $\bar{G}_{\ell}\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)})}{\partial x_{\ell}}\frac{\partial f(x^{(\alpha)})}{\partial x_{\ell}}^{\top}.$
  3. 3.
    Neural feature matrix (NFM). Given a neural network $f$ with a weight matrix $W^{(\ell)}$ at layer $\ell$, the NFM $F_{\ell}$ is defined as,
    $F_{\ell}\equiv(W^{(\ell)})^{\top}W^{(\ell)}.$
  4. 4.
    Neural feature correlation (NFC). Defined with respect to a layer of a neural network, this is the correlation $\rho\left(F_{\ell},\bar{G}_{\ell}\right).$
  5. 5.
    Neural feature ansatz (NFA). This refers to the statement that the NFC will have correlation approximately equal to $1$.
  6. 6.
    Pre-activation tangent kernel (PTK). This is the kernel function corresponding, defined with respect to a neural network and a layer $\ell$, that evaluates two inputs $x,z$ to $\frac{\partial f(x)}{\partial h_{\ell}}\cdot\frac{\partial f(z)}{\partial h_{\ell}}\in\mathbb{R}$.
  7. 7.
    PTK feature covariance. For a given layer of a neural network and a dataset $X$, the PTK feature covariance $K^{(\ell)}$ is defined as $K^{(\ell)}\equiv\frac{\partial f(X)}{\partial h_{\ell}}^{\top}\frac{\partial f(X)}{\partial h_{\ell}}\in\mathbb{R}^{k_{\ell}\times k_{\ell}}$.
  8. 8.
    Centered NFC (C-NFC). Consider a layer $W$ of a neural network with PTK feature covariance $K$ with respect to that layer. Let $\bar{W}\equiv W-W_{0}$, where $W_{0}$ are weights at initialization. Then, the C-NFC is equivalent to the correlation $\rho\left(\bar{W}^{\top}\bar{W},\bar{W}^{\top}K\bar{W}\right)$.
  9. 9.
    Uncentered NFC (UC-NFC). This quantity is identical to the NFC defined above.
Appendix B Connection of the PTK to the empirical NTK
We note the following connection between the PTK entries and the empirical NTK (ENTK). In particular, the PTK is a significant component of the ENTK.
Proposition 4 (Pre-activation to neural tangent identity).
Consider a depth-$L$ neural network $f(x)$ with inputs $X_{\ell}\in\mathbb{R}^{n\times k_{\ell}}$ to weight matrices $W^{(\ell)}\in\mathbb{R}^{k_{\ell}\times k_{\ell}}$ for $\ell\in[L]$. Consider the empirical NTK where gradients are taken only with respect to $W^{(\ell)}$ evaluated between two points $x$ and $z$ denoted by $\hat{\Theta}_{\ell}(x_{0},z_{0})=\langle\frac{\partial f(x_{0})}{\partial W^{(\ell)}},\frac{\partial f(z_{0})}{\partial W^{(\ell)}}\rangle$. Then for all $\ell$, $\hat{\Theta}_{\ell}$ satisfies,
$\hat{\Theta}_{\ell}(x_{0},z_{0})=\mathcal{K}^{(\ell)}(x_{0},z_{0})\cdot x_{\ell}^{\top}z_{\ell}~{}.$
Proof of Proposition 4.
Note that $\frac{\partial f(x_{0})}{\partial W^{(\ell)}}=\frac{\partial f(x_{0})}{\partial h_{\ell}}x_{0}^{\top}\in\mathbb{R}^{k_{\ell}\times k_{\ell}}$. Then,
$\hat{\Theta}_{\ell}(x_{0},z_{0})$ $=\text{tr}\left(\frac{\partial f(x_{0})}{\partial h_{\ell}}x_{\ell}^{\top}z_{\ell}\frac{\partial f(z_{0})}{\partial h_{\ell}}^{\top}\right)$
$=\frac{\partial f(x_{0})}{\partial h_{\ell}}^{\top}\frac{\partial f(z_{0})}{\partial h_{\ell}}\cdot x_{\ell}^{\top}z_{\ell}$
$=\mathcal{K}^{(\ell)}(x_{0},z_{0})\cdot x_{\ell}^{\top}z_{\ell}~{}.$
Appendix C Additional centerings of the NFC
Double-centered NFC
One may additionally center the PTK feature map to understand the co-evolution of the PTK feature covariance and the weight matrices. We consider such a centering that we refer to as the double-centered NFC, where we measure $\rho\left((\bar{W}^{(\ell)})^{\top}\bar{W}^{(\ell)},(\bar{W}^{(\ell)})^{\top}\bar{K}^{(\ell)}\bar{W}^{(\ell)}\right)$, where $\bar{K}^{(\ell)}=\left(\frac{\partial f(X)}{\partial h_{\ell}}-\frac{\partial f_{0}(X)}{\partial h_{\ell}}\right)^{\top}\left(\frac{\partial f(X)}{\partial h_{\ell}}-\frac{\partial f_{0}(X)}{\partial h_{\ell}}\right)$, and $f_{0}$ is the neural network at initialization.
However, the double-centered NFC term corresponds to higher-order dynamics that do not significantly contribute the centered and uncentered NFC (Figure 5) when initialization is large or for early periods of training. Note however this term becomes relevant over longer periods of training.
Figure 5: Ratio of the unnormalized double-centered NFC to the centered NFC throughout neural network training. In particular, we plot $\text{tr}\left(\bar{W}^{\top}\bar{W}\bar{W}^{\top}\bar{K}\bar{W}\right)\cdot{\text{tr}\left(\bar{W}^{\top}\bar{W}\bar{W}^{\top}K\bar{W}\right)}^{-1}$ throughout training for both layers of a two-hidden layer MLP with ReLU activations.
Isolating alignment of the PTK to the initial weight matrix
One may also center just the PTK feature map, while substituting the initial weights for $W$ to isolate how the PTK feature covariance aligns to the weight matrices. To measure this alignment, we consider the PTK-centered NFC, which is defined as the correlation $\rho\left((W_{0}^{(\ell)})^{\top}W_{0}^{(\ell)},(W_{0}^{(\ell)})^{\top}\bar{K}^{(\ell)}W_{0}^{(\ell)}\right)~{}$, where $W_{0}^{(\ell)}$ is the initial weight matrix at layer $\ell$.
However, this correlation decreases through training, indicating that the correlation of these quantities does not drive alignment between the uncentered NFM and AGOP (Figure 6).
(a) Isotropic data.
(b) Data spectrum with decay $\lambda_{k}\sim\frac{1}{1+k^{2}}$.
Figure 6: PTK-centered NFC throughout training for both layers of a two-hidden layer MLP with ReLU activations on Gaussian data with two different spectra.
Appendix D Extending our theoretical predictions to depth and general activations
Precise predictions of the C-NFC become more complicated with additional depth and general activation functions. However, we note that the deep C-NFC will remain sensitive to a first-order approximation in which $K$ is replaced by its expectation. We demonstrate that this term qualitatively captures the behavior of the C-NFC for 2 hidden layer architectures with quadratic and, to a lesser extent, ReLU activation functions in Figure 7. In this experiment, we sample Gaussian data with mean 00 and covariance with a random eigenbasis. We parameterize the eigenvalue decay of the covariance matrix by a parameter $\alpha$, called the data decay rate, so that the eigenvalues have values $\lambda_{k}=\frac{1}{1+k^{\alpha}}$. As $\alpha$ approaches 00 or $\infty$ the data covariance approaches a projector matrix.
In this experiment, we see that the data covariance spectrum will also parameterize the eigenvalue decay of $\mathbb{E}\left[K\right]$, allowing us to vary how close the expected PTK matrix (and its dual, the PTK feature covariance) is from a projector, where the NFA holds exactly. We see that for intermediate values of $\alpha$, both the observed and the predicted derivatives of the C-NFC decreases in value.
We plot the observed values in two settings corresponding to different asymptotic regimes. One setting is the proportional regime, where $n=k=d=128$. The other is the NTK regime where $n=d=128$ and $k=1024$. For the quadratic case, as the network approaches infinite width, the prediction more closely matches the observed values. Additional terms corresponding to the nonlinear part of $\phi^{\prime}$ in ReLU networks, the derivative of the activation function, are required to capture the correlation more accurately in this case.
(a) Quadratic.
(b) ReLU.
Figure 7: Observed versus the first-order predicted C-NFC for the input to the first layer of a two hidden layer MLP. The dashed line is neural network width $k=n=d=128$, where $n$ and $d$ are the number of data point and data dimension, respectively, while the solid line uses $n=d=128$ and $k=1024$.
Appendix E Free probability calculations of C-NFC
In order to understand the development of the NFC, we analyze the centered NFC in the limit that learning rate is much smaller than the initialization for a one hidden layer MLP with quadratic activations. We write this particular network as,
$f(x)=a^{\top}(Wx)^{2}~{},$
where $a\in\mathbb{R}^{1\times k}$ and $W\in\mathbb{R}^{k\times d}$, where $d$ is the input dimension and $k$ is the width. In this case, the NFC has the following form,
$\rho\left(F,\bar{G}\right)=\frac{\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)}{\text{tr}\left((X^{\top}Y\mathcal{K}YX)^{2}\right)^{-1/2}\text{tr}\left((X^{\top}Y\mathcal{K}^{2}YX)^{2}\right)^{-1/2}}~{},$ (8)
where $\mathcal{K}=XW^{\top}{\rm diag}\left(a\right)^{2}WX^{\top}$.
We assume three properties hold in the finite dimensional case we consider, that will hold asymptotically in the infinite dimensional limit.
Assumption 5 (Self-averaging).
We assume that computing the average of the NFC quantities across initializations is equal to the quantities themselves in the high-dimensional limit.
Assumption 6 (Asymptotic freeness).
We assume that the collections $\{X,Y\}$ and $\{W,a\}$ are asymptotically free with respect to the operator $\mathbb{E}\left[{\rm tr}(\cdot)\right]$, where ${\rm tr}[M]=\frac{1}{n}\sum_{i=1}^{n}M_{ii}$.
Assumption 7 (Commutativity of expectation).
We will also make the approximation that the expectation commutes with ratio and square root.
$\mathbb{E}\left[\rho(A,B)\right]=\mathbb{E}\left[\frac{{\rm tr}(A^{\top}B)}{\sqrt{{\rm tr}(A^{\top}A){\rm tr}(B^{\top}B)}}\right]\approx\frac{\mathbb{E}\left[{\rm tr}(A^{\top}B)\right]}{\sqrt{\mathbb{E}\left[{\rm tr}(A^{\top}A)]{\rm E}[{\rm tr}(B^{\top}B)\right]}}$ (9)
We will compute the expected values of the centered NFC under these assumptions. In the remainder of the section we will drop the $\mathbb{E}\left[\cdot\right]$ in the trace for ease of notation.
E.1 Free probability identities
The following lemmas will be useful: let $\{\bar{C}_{i}\}$ and $\{R_{i}\}$ be freely independent of each other with respect to ${\rm tr}$, with ${\rm tr}[\bar{C}_{i}]=0$. Alternating words have the following products:
${\rm tr}[\bar{C}_{1}R_{1}]=0$ (10)
${\rm tr}[\bar{C}_{1}R_{1}\bar{C}_{2}R_{2}]={\rm tr}[R_{1}]{\rm tr}[R_{2}]{\rm tr}[\bar{C}_{1}\bar{C}_{2}]$ (11)
${\rm tr}[\bar{C}_{1}R_{1}\bar{C}_{2}R_{2}\bar{C}_{3}R_{3}]={\rm tr}[R_{1}]{\rm tr}[R_{2}]{\rm tr}[R_{3}]{\rm tr}[\bar{C}_{1}\bar{C}_{2}\bar{C}_{3}]$ (12)
$\begin{split}&{\rm tr}[\bar{C}_{1}R_{1}\bar{C}_{2}R_{2}\bar{C}_{3}R_{3}\bar{C}_{4}R_{4}]={\rm tr}[R_{1}]{\rm tr}[R_{2}]{\rm tr}[R_{3}]{\rm tr}[R_{4}]{\rm tr}[\bar{C}_{1}\bar{C}_{2}\bar{C}_{3}\bar{C}_{4}]+\\ &{\rm tr}[R_{1}]{\rm tr}[R_{3}]{\rm tr}[\bar{R}_{2}\bar{R}_{4}]{\rm tr}[\bar{C}_{1}\bar{C}_{2}]{\rm tr}[\bar{C}_{3}\bar{C}_{4}]+{\rm tr}[R_{2}]{\rm tr}[R_{4}]{\rm tr}[\bar{R}_{1}\bar{R}_{3}]{\rm tr}[\bar{C}_{2}\bar{C}_{3}]{\rm tr}[\bar{C}_{1}\bar{C}_{4}]\end{split}$ (13)
where $\bar{R}_{i}\equiv R_{i}-{\rm tr}[R_{i}]$.
Applying these identities to the one hidden layer quadratic case, we use the following definitions:
$R=W^{\top}{\rm diag}\left(a^{2}\right)W,~{}A=(X^{\top}YX)^{2},~{}B=X^{\top}X$ (14)
Crucially, $R$ is freely independent of the set $\{A,B\}$. We will also use the notation $\bar{M}$ to indicated the centered version of $M$, $\bar{M}=M-{\rm tr}[M]$.
E.2 Numerator term of NFC
The numerator in Equation (8) is
$\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)=\text{tr}\left(ARARBR\right)$ (15)
Re-writing $A=\bar{A}+{\rm tr}[A]$ and $B=\bar{B}+{\rm tr}[B]$ we have:
$\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)=\text{tr}\left((\bar{A}+{\rm tr}[A])R(\bar{A}+{\rm tr}[A])R(\bar{B}+{\rm tr}[B])R\right)$ (16)
This expands to
$\begin{split}\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)&=\text{tr}\left(\bar{A}R\bar{A}R\bar{B}R\right)+2\text{tr}\left(A\right)\text{tr}\left(\bar{A}R\bar{B}R^{2}\right)+\text{tr}\left(B\right)\text{tr}\left(\bar{A}R\bar{A}R^{2}\right)\\ &\text{tr}\left(A\right)^{2}\text{tr}\left(\bar{B}R^{3}\right)+2\text{tr}\left(A\right)\text{tr}\left(B\right)\text{tr}\left(\bar{A}R^{3}\right)+\text{tr}\left(A\right)^{2}\text{tr}\left(B\right)\text{tr}\left(R\right)^{3}\end{split}$ (17)
Using the identities we arrive at:
$\begin{split}\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)&={\rm tr}[A]^{2}{\rm tr}[B]\text{tr}\left(R^{3}\right)\\ &+2{\rm tr}[A]{\rm tr}[R^{2}]{\rm tr}[R]\text{tr}\left(\bar{A}\bar{B}\right)\\ &+{\rm tr}[B]{\rm tr}[R]{\rm tr}[R^{2}]\text{tr}\left(\bar{A}^{2}\right)\\ &+{\rm tr}[R]^{3}\text{tr}\left(\bar{A}^{2}\bar{B}\right)\end{split}$ (18)
E.3 First denominator term of NFC
The first denominator term in Equation (8) is
$\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}YX\right)=\text{tr}\left(ARAR\right)$ (19)
This is a classic free probability product:
$\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}YX\right)={\rm tr}[A^{2}]{\rm tr}[R]^{2}+{\rm tr}[A]^{2}\text{tr}\left(R^{2}\right)-{\rm tr}[A]^{2}{\rm tr}[R]^{2}$ (20)
which can be derived from the lemmas.
E.4 Second denominator term of NFC
For the second denominator term of Equation (8) we have
$\text{tr}\left(X^{\top}Y\mathcal{K}^{2}YXX^{\top}Y\mathcal{K}^{2}YX\right)=\text{tr}\left(ARBRARBR\right)$ (21)
Expanding the first $A$ we have
$\text{tr}\left(X^{\top}Y\mathcal{K}^{2}YXX^{\top}Y\mathcal{K}^{2}YX\right)=\text{tr}\left(\bar{A}RBRARBR\right)+{\rm tr}[A]\text{tr}\left(R^{2}BRARB\right)$ (22)
Next we expand the first $B$:
$\begin{split}\text{tr}\left(X^{\top}Y\mathcal{K}^{2}YXX^{\top}Y\mathcal{K}^{2}YX\right)&=\text{tr}\left(\bar{A}R\bar{B}RARBR\right)+{\rm tr}[B]\text{tr}\left(\bar{A}R^{2}ARBR\right)\\ &+{\rm tr}[A]{\rm tr}[B]\text{tr}\left(R^{3}ARB\right)+{\rm tr}[A]\text{tr}\left(R^{2}\bar{B}RARB\right)\end{split}$ (23)
The next $A$ gives us
$\begin{split}\text{tr}\left(X^{\top}Y\mathcal{K}^{2}YXX^{\top}Y\mathcal{K}^{2}YX\right)&=\text{tr}\left(\bar{A}R\bar{B}R\bar{A}RBR\right)+2{\rm tr}[A]\text{tr}\left(\bar{A}R\bar{B}R^{2}BR\right)+{\rm tr}[B]\text{tr}\left(\bar{A}R^{2}\bar{A}RBR\right)\\ &+2{\rm tr}[A]{\rm tr}[B]\text{tr}\left(R^{3}\bar{A}RB\right)+{\rm tr}[A]^{2}{\rm tr}[B]\text{tr}\left(R^{4}B\right)+{\rm tr}[A]^{2}\text{tr}\left(R^{2}\bar{B}R^{2}B\right)\end{split}$ (24)
Expanding the final $B$ we have
$\begin{split}\text{tr}\left(X^{\top}Y\mathcal{K}^{2}YXX^{\top}Y\mathcal{K}^{2}YX\right)&=\text{tr}\left(\bar{A}R\bar{B}R\bar{A}R\bar{B}R\right)+2{\rm tr}[B]\text{tr}\left(\bar{A}R\bar{B}R\bar{A}R^{2}\right)+2{\rm tr}[A]\text{tr}\left(\bar{A}R\bar{B}R^{2}\bar{B}R\right)\\ &+4{\rm tr}[A]{\rm tr}[B]\text{tr}\left(R^{3}\bar{A}R\bar{B}\right)+2{\rm tr}[A]{\rm tr}[B]^{2}\text{tr}\left(R^{4}\bar{A}\right)+2{\rm tr}[A]^{2}{\rm tr}[B]\text{tr}\left(R^{4}\bar{B}\right)\\ &+{\rm tr}[A]^{2}{\rm tr}[B]^{2}{\rm tr}[R^{4}]+{\rm tr}[A]^{2}{\rm tr}[R^{2}\bar{B}R^{2}\bar{B}]+{\rm tr}[B]^{2}{\rm tr}[R^{2}\bar{A}R^{2}\bar{A}]\end{split}$ (25)
Now all terms are in the form of alternating products from the lemma. This means we can factor out the non-zero traces of the other terms. Simplifying we have:
$\begin{split}\text{tr}\left(X^{\top}Y\mathcal{K}^{2}YXX^{\top}Y\mathcal{K}^{2}YX\right)&={\rm tr}[R]^{4}\text{tr}\left((\bar{A}\bar{B})^{2}\right)+2{\rm tr}[R]^{2}({\rm tr}[R^{2}]-{\rm tr}[R]^{2}){\rm tr}[\bar{A}\bar{B}]^{2}\\ &+2{\rm tr}[R]^{2}{\rm tr}[R^{2}]\left({\rm tr}[B]\text{tr}\left(\bar{A}^{2}\bar{B}\right)+{\rm tr}[A]\text{tr}\left(\bar{A}\bar{B}^{2}\right)\right)\\ &+4{\rm tr}[A]{\rm tr}[B]{\rm tr}[R^{3}]{\rm tr}[R]\text{tr}\left(\bar{A}\bar{B}\right)+{\rm tr}[A]^{2}{\rm tr}[B]^{2}{\rm tr}[R^{4}]\\ &+{\rm tr}[A]^{2}{\rm tr}[\bar{B}^{2}]{\rm tr}[R^{2}]^{2}+{\rm tr}[\bar{A}]^{2}{\rm tr}[B^{2}]{\rm tr}[R^{2}]^{2}\end{split}$ (26)
All terms of the NFC are now in terms of traces of the matrices $A$, $B$, and $R$ and functions on each term separately. The matrices $A$ and $B$ are determined by the data, while the moments of the eigenvalues of $R$ are determined by the initialization distribution of the weights in the neural network, and neither training nor the data.
Appendix F Alignment reversing dataset
The data consists of a mixture of two distributions from which two subsets of the data $X^{(1)}$ and $X^{(2)}$ are sampled from, and is parametrized by a balance parameter $\gamma\in(0,1]$ and two variance parameters $\epsilon_{1},\epsilon_{2}>0$. The subset $X^{(1)}$ which has label $y_{1}=1$ and constitutes a $\gamma$ fraction of the entire dataset, is sampled from a multivariate Gaussian distribution with mean 00 and covariance
$\Sigma=\bm{1}\bm{1}^{\top}+\epsilon_{1}\cdot I~{}.$
We construct $X^{(2)}$ by the following procedure:
  1. 1.
    Extract singular values $S_{1}$ and right singular vectors $U_{1}$ from a singular-value decomposition (SVD) of $(X^{(1)})^{\top}X^{(1)}$.
  2. 2.
    Extract the left singular vectors $V_{2}$ from a sample $\tilde{X}_{2}$ that is sampled from the same distribution as $X^{(1)}$.
  3. 3.
    Construct $X^{(2)}=V_{2}S_{1}^{-1}U_{1}^{\top}$.
  4. 4.
    Where $X=X^{(1)}\oplus X^{(2)}$, Set $X\leftarrow X+\epsilon_{2}Z$, where $Z\sim\mathcal{N}(0,I)$.
  5. 5.
    Set $y\leftarrow y+10^{-5}\cdot\bm{1}$.
Note that $U_{1}S_{1}^{-1}V_{2}^{\top}V_{2}S_{1}^{-1}U_{1}^{\top}=U_{1}S_{1}^{-2}U_{1}^{\top}=((X^{(1)})^{\top}X^{(1)})^{-2}$, therefore, we should set $X^{(2)}=V_{2}S_{1}^{-1}U_{1}^{\top}$ to get $(X^{(2)})^{\top}X^{(2)}=((X^{(1)})^{\top}X^{(1)})^{-2}$. Regarding the variance parameters, in practice we set $\epsilon_{1}=0.5$ and $\epsilon_{2}=10^{-2}$.
Proposition 8 (Expected NFM and AGOP).
For a one hidden layer quadratic network, $f(x)=a^{\top}(Wx)^{2}$, with $a\sim\mathcal{N}(0,I)$ and $W\sim\frac{1}{\sqrt{k}}\cdot\mathcal{N}(0,I)$,
$\mathbb{E}_{a,W}\left[\dot{W}^{\top}\dot{W}\right]=(X^{\top}YX)^{2}~{},$
and,
$\mathbb{E}_{a,W}\left[\dot{W}^{\top}K\dot{W}\right]$ $=3\cdot\text{tr}\left(X^{\top}X\right)\cdot(X^{\top}YX)^{2}$
$+6X^{\top}YXX^{\top}XX^{\top}YX$
Proof of Proposition 8.
$\mathbb{E}\left[\dot{W}^{\top}\dot{W}\right]$ $=X^{\top}YX\mathbb{E}\left[W_{0}^{\top}{\rm diag}\left(a\right)^{2}W_{0}\right]X^{\top}YX$
$=(X^{\top}YX)^{2}~{}.$
Further,
$\mathbb{E}\left[\dot{W}^{\top}K\dot{W}\right]$ $=X^{\top}Y\mathbb{E}\left[K^{2}\right]YX~{}.$
We note that,
$K^{2}$ $=W_{0}^{\top}{\rm diag}\left(a\right)^{2}W_{0}X^{\top}XW_{0}^{\top}{\rm diag}\left(a\right)^{2}W_{0}$ (27)
$=\sum_{s_{1},s_{2}}^{k}\sum_{\alpha}^{n}\sum_{p_{1},p_{2}}^{d}$ (28)
$a_{s_{1}}^{2}a_{s_{2}}^{2}W_{s_{1},p_{1}}X_{\alpha,p_{1}}X_{\alpha,p_{2}}W_{s_{2},p_{2}}XW_{s_{1}}W_{s_{2}}^{\top}X^{\top}~{}.$ (29)
Therefore, applying Wick’s theorem, element $i,j$ of $K^{2}$ satisfies,
$\mathbb{E}\left[K^{2}_{ij}\right]$ $=\sum_{s}^{k}\sum_{\alpha}^{n}\sum_{p_{1},p_{2}}^{d}\mathbb{E}\left[a_{s}^{4}W_{s,p_{1}}X_{\alpha,p_{1}}X_{\alpha,p_{2}}W_{s,p_{2}}X_{i}^{\top}W_{s}W_{s}^{\top}X_{j}\right]$
$=\sum_{s}^{k}\sum_{\alpha}^{n}\sum_{p_{1},p_{2},q_{1},q_{2}}^{d}$
$\mathbb{E}\left[a_{s}^{4}W_{s,p_{1}}W_{s,p_{2}}W_{s,q_{1}}W_{s,q_{2}}X_{\alpha,p_{1}}X_{\alpha,p_{2}}X_{i,q_{1}}X_{j,q_{2}}\right]$
$=3\sum_{s}^{k}\sum_{\alpha}^{n}\sum_{p_{1},p_{2},q_{1},q_{2}}^{d}$
$\Big{(}\mathbb{E}\left[W_{s,p_{1}}W_{s,p_{2}}\right]\mathbb{E}\left[W_{s,q_{1}}W_{s,q_{2}}\right]+$
$\mathbb{E}\left[W_{s,p_{1}}W_{s,q_{1}}\right]\mathbb{E}\left[W_{s,p_{2}}W_{s,p_{2}}\right]+$
$\mathbb{E}\left[W_{s,p_{1}}W_{s,q_{2}}\right]\mathbb{E}\left[W_{s,p_{2}}W_{s,q_{1}}\right]\Big{)}$
$\cdot X_{\alpha,p_{1}}X_{\alpha,p_{2}}X_{i,q_{1}}X_{j,q_{2}}$
$=3\sum_{\alpha}^{n}\sum_{p_{1},p_{2},q_{1},q_{2}}^{d}$
$\Big{(}\delta_{p_{1}p_{2}}\delta_{q_{1}q_{2}}+\delta_{p_{1}q_{1}}\delta_{p_{2}q_{2}}+\delta_{p_{1}q_{2}}\delta_{p_{2}q_{1}}\Big{)}$
$\cdot X_{\alpha,p_{1}}X_{\alpha,p_{2}}X_{i,q_{1}}X_{j,q_{2}}$
$=3\sum_{\alpha}^{n}\Big{(}\sum_{p_{1},q_{1}}^{d}X_{\alpha,p_{1}}X_{\alpha,p_{1}}X_{i,q_{1}}X_{j,q_{1}}$
$+\sum_{p_{1},p_{2}}^{d}X_{\alpha,p_{1}}X_{\alpha,p_{2}}X_{i,p_{1}}X_{j,p_{2}}$
$+\sum_{p_{1},p_{2}}^{d}X_{\alpha,p_{1}}X_{\alpha,p_{2}}X_{i,p_{2}}X_{j,p_{1}}\Big{)}$
$=3\cdot\text{tr}\left(X^{\top}X\right)\cdot X_{i}^{\top}X_{j}+3\sum_{\alpha}^{n}X_{\alpha}^{\top}X_{i}X_{\alpha}^{\top}X_{j}$
$+3\sum_{\alpha}^{n}X_{\alpha}^{\top}X_{j}X_{\alpha}^{\top}X_{i}$
$=3\cdot\text{tr}\left(X^{\top}X\right)\cdot X_{i}^{\top}X_{j}+3X_{i}X^{\top}XX_{j}+3X_{j}X^{\top}XX_{i}~{}.$
Finally, we conclude,
$\mathbb{E}\left[K^{2}\right]=3\left(\text{tr}\left(X^{\top}X\right)XX^{\top}+2XX^{\top}XX^{\top}\right)~{},$
giving the second statement of the proposition. ∎
Appendix G Varying the data distribution
We verify that our observations for isotropic Gaussian data hold even when the data covariance has a significant spectral decay. (Figures 9 and 10). We again consider Gaussian data that is mean 00 and where the covariance is constructed from a random eigenbasis. In Figure 9, we substitute the eignevalue decay as $\lambda_{k}\sim\frac{1}{1+k}$, while in Figure 10, we use $\lambda_{k}\sim\frac{1}{1+k^{2}}$. We plot the values of the UC-NFC, C-NFC, train loss, and test loss throughout training for the first and second layer of a two hidden layer network with ReLU activations, while additionally varying initialization scale. Similar to Figure 8, we observe that the C-NFC is more robust to the initialization scale than the UC-NFC, and UC-NFC value become high through training, while being small at initialization. We see that the test loss improves for smaller initializations, where the value of the C-NFC and UC-NFC are higher.
Figure 8: Centered neural feature correlations. Data covariance decay rate $\lambda_{k}=1$. Top row is layer 1, bottom row is layer 2. Train (test) losses are scaled by the maximum train (test) loss achieved so that they are between 00 and $1$.
Figure 9: Centered neural feature correlations. Data covariance decay rate $\lambda_{k}\sim\frac{1}{1+k}$. Top row is layer 1, bottom row is layer 2. Train (test) losses are scaled by the maximum train (test) loss achieved so that they are between 00 and $1$.
Figure 10: Centered neural feature correlations. Data covariance decay rate $\lambda_{k}\sim\frac{1}{1+k^{2}}$. Top row is layer 1, bottom row is layer 2. Train (test) losses are scaled by the maximum train (test) loss achieved so that they are between 00 and $1$.
Appendix H Effect of initialization on feature learning
We see that when initialization is small, the C-NFC and UC-NFC are high at the end of training with and without fixing the learning speed (Figure 4). This is reflected by the quality of the features learned by the NFM and the qualitative similarity of the NFM and AGOP at small initialization scale (Figure 11). Further, we notice that as we increase initialization, without fixing speeds, the correspondence between the NFM and decreases and the quality of the NFM features decreases (at a faster rate than the AGOP). Strikingly, when learning speeds are fixed, the quality of the features in the AGOP and NFM becomes invariant to the initialization scale.
(a) NFM, speeds not fixed.
(b) AGOP, speeds not fixed.
(c) NFM, speeds fixed.
(d) AGOP, speeds fixed.
Figure 11: The NFM at the end of training as a function of initialization scale in the first layer weights, with and without fixing learning speeds. The task again is the chain monomial with rank $r=5$. The title of each plot is the initialization scale of the first layer $s_{0}$.
Appendix I Experimental details
We describe the neural network training and architectural hyperparameters in the experiments of this paper. Biases were not used for any networks. Further, in all polynomial tasks, we scaled the label vector to have standard deviation $1$.
Corrupted AGOP
For the experiments in Figure 1, we used $n=384$ data points, $d=32$, $k=128$ as the width in all layers, isotropic Gaussian data, initialization scale $0.01$ in the first layer and default scale in the second. We used ReLU activations and two hidden layers. For the experiments in Figure 8,9,10,5, and 6, we used a two hidden layer network with ReLU activations, learning rate $0.05$, 800 steps of gradient decent, and took correlation/covariance measurements every 5 steps.
C/UC-NFC calculations on real datasets
We describe the experimental details for Figure 2. For (A,B) we trained a five layer MLP on the first 50,000 datapoints of Streetview House Numbers (SVHN), CIFAR-10, CIFAR-100, STL-10, MNIST, and German Traffic Sign Recognition Benchmark (GTSRB) datasets. We used the default PyTorch initialization (scale of $1$) for all layers. We used width 256 in all layers, and trained with SGD with batch size 128. For SVHN and CIFAR, we trained for 150 epochs with learning rate 0.2. For STL-10 and GTSRB we used learning rate 0.1 for 150 epochs. For MNIST we trained for 50 epochs with learning rate 10. For the VGG-11 experiments on CIFAR-10, we used the default architecture from torchvision with batch-norm layers removed. We trained for 50 epochs and learning rate 1. For experiments with a GPT-family model, we adapted the model and dataset from NanoGPT (https://github.com/karpathy/nanoGPT). We used all the default settings and the default Adam optimizer, but used no weight decay, learning rate 5e-3, and removed all dropout layers. We also reduced the number of attention layers to 3 from the original 6. We trained on the Shakespeare characters dataset.
Alignment reversing dataset
For the experiments in Figure 3, we used $k=n=d=1024$ for the width, dataset size, and input dimension, respectively. Further, the traces of powers of $F_{a}$ are averaged over 30 neural net seeds to decouple these calculated values from the individual neural net seeds. The mean value plotted in the first two squares of figure is computed over 10 data seeds.
SLO experiments
For the SLO figures (Figures 11, 4), we use isotropic Gaussian data, 600 steps of gradient descent. The learning rates are chosen based on initialization scale in the first layer. For initialization scales 1, 0.1, 0.01, and 0.001, we used learning rates $0.03$, $0.1$, $0.2$, $0.4$, respectively. We again used two hidden layers with ReLU activations. We chose $n=256$, $d=32$, and $k=256$ as the width. We divided the linear readout weights by $0.01$ at initialization to promote feature learning, and modified SLO to scale gradients by $(\epsilon+\|\nabla\mathcal{L}\|)^{-1}$, rather than just the inverse of the norm of the gradient, for $\epsilon=0.1$. This technique smooths the training dynamics as the parameters approach a loss minimum, allowing the network to interpolate the labels.
Predictions with depth
For the Deep C-NFC predictions (Figure 7), we used $n=128$, $d=128$, initialization scale of $1$. The low rank task is just the chain monomial of rank $r=5$. The high rank polynomial task is $y(x)=\sum_{i=1}^{d}(Qx)_{i}^{2}$, where $Q\in\mathbb{R}^{d\times d}$ is a matrix with standard normal entries.
Figures 16 and 17
For the experiments on the SVHN dataset, we train a four hidden layer neural network with ReLU activations, initialization scale $1.0$ in all layers and width $256$. For SVHN, we subset the dataset to 4000 points. We train for 3000 epochs with learning rate 0.2 for standard training, and 0.3 for SLO, and take NFC measurements every 50 epochs. For SLO, we set $C_{0}=2.5$, $C_{1}=C_{2}=0.4$, and relaxation parameter $\epsilon=0.2$. We pre-process the dataset so that each pixel is mean 00 and standard deviation $1$. For the experiments on CelebA, we train a two hidden layer network on a balanced subset of 7500 points with Adam with learning rate $0.0001$ and no weight decay. We use initialization scale 0.02 in the first layer, and width 128. We train for 500 epochs. We pre-process the dataset by scaling the pixel values to be between 0 and 1.
Code availability
We make the code available for the experiments in Figure 2 available at https://anonymous.4open.science/r/centered_NFA-4795/ (for MLP+VGG).
Appendix J Additional SLO experiments
We demonstrate the SLO can be applied adaptively to increase the strength of the UC-NFC in all layers of a deep network on the chain monomial task of rank $r=3$. We train a three hidden layer MLP with ReLU activations and an initialization scale of $0.1$ by SLO, and find that all layers finish at the same high UC-NFC (Figure 12). Further, this final UC-NFC value is higher than the highest UC-NFC achieved by any layer with standard training. The generalization loss is also lower with SLO on this example, corresponding to better feature learning (through the UC-NFC).
Figure 12: Training with SLO where the learning speeds are chosen adaptively based on the UC-NFC values of all layers. The dashed lines correspond to training with standard GD.
At every time step we choose $C_{i}=s$ for the layer $i$ with the smallest UC-NFC correlation value, while setting $C_{j}=s^{-1}$ for all other layers, with $s=20$. We again modify SLO by dividing the gradients by ϵitalic-$\epsilon\|\nabla\mathcal{L}\|$ for $\epsilon=0.01$. The learning rate is set to $0.05$ in SLO and $0.25$ for the standard training (gradient descent), and the networks are trained for 500 epochs. We sample $n=256$ points with $d=32$, and use width $k=256$.
Appendix K Additional C-NFC/UC-NFC plots across architectures
Figure 13: Full trajectories for the C/UC-NFC of a five hidden layer MLP trained on six datasets, averaged over three seeds, with large ($1.0$) initialization scale. The blue curves are the UC-NFC, while green curves are the C-NFC.
Figure 14: Losses and normalized changes in weights across datasets for a five hidden layer MLP. The change in weight is measured as $\|W-W_{0}\|\|W\|^{-1}$. First column of all subfigures are the losses, while the second columns are the weight changes.
Figure 15: Losses and normalized changes in weights across datasets for VGG-11 on CIFAR-10 and GPT on Shakespeare character generation. The change in weight is measured as $\|W-W_{0}\|\|W\|^{-1}$.
First column of (A) and (B) are the losses, while the remaining columns are the weight changes.
Appendix L Additional experiments on real datasets
We replicate Figures 1 and 4 on celebrity faces (CelebA) and Street View House Numbers (SVHN). We begin by showing that one can disrupt the NFC correspondence by replacing the PTK feature covariance with a random matrix of the same spectral decay. For this example, we measure the Pearson correlation, which subtracts the mean of the image. I.e. $\bar{\rho}(A,B)\equiv\rho\left(A-m(A),B-m(B)\right)$, where $m(A),m(B)$ are the average of the elements of $A$ and $B$.
Figure 16: Various feature learning measures for the CelebA binary subtask of predicting glasses. The diagonals of the NFM $\left(W^{\top}W\right)$ (first plot) and AGOP $\left(W^{\top}KW\right)$ (second plot) of a fully-connected network are similar to each other. Replacing $K$ with a symmetric matrix $Q$ with the same spectrum but independent eigenvectors obscures the low rank structure (third plot), and reduces the Pearson correlation of the diagonal from $\bar{\rho}\left({\rm diag}\left(F\right),{\rm diag}\left(\bar{G}\right)\right)=0.91$ to $\bar{\rho}\left({\rm diag}\left(F\right),{\rm diag}\left(W^{\top}QW\right)\right)=0.04$.
Figure 17: We demonstrate on the SVHN dataset, with a 4 hidden layer neural network with large initialization scale, how SLO can improve the strength of the UC-NFC, the C-NFC, the ratio of the unnormalized C-NFC to UC-NFC (plot (a)) and the feature quality (plots (b) and (c)). In plots (b) and (c), we visualize the diagonal of the NFM and AGOP for the first layer of the trained network, where SLO was applied with $C_{0}=2.5$, $C_{1}=C_{2}=0.4$.